# SANeRF-HQ: Segment Anything for NeRF in High Quality

### 03 Dec 2023 v1

@@ Abstract
Recently, the Segment Anything Model (SAM) has showcased remarkable capabilities of zero-shot segmentation, while NeRF (Neural Radiance Fields) has gained popularity as a method for various 3D problems beyond novel view synthesis. Though there exist initial attempts to incorporate these two methods into 3D segmentation, they face the challenge of accurately and consistently segmenting objects in complex scenarios. In this paper, we introduce the Segment Anything for NeRF in High Quality (SANeRF-HQ) to achieve high quality 3D segmentation of any object in a given scene. SANeRF-HQ utilizes SAM for open-world object segmentation guided by user-supplied prompts, while leveraging NeRF to aggregate information from different viewpoints. To overcome the aforementioned challenges, we employ density field and RGB similarity to enhance the accuracy of segmentation boundary during the aggregation. Emphasizing on segmentation accuracy, we evaluate our method quantitatively on multiple NeRF datasets where high-quality ground-truths are available or manually annotated. SANeRF-HQ shows a significant quality improvement over previous state-of-the-art methods in NeRF object segmentation, provides higher flexibility for object localization, and enables more consistent object segmentation across multiple views. Additional information can be found at https://lyclyc52.github.io/SANeRF-HQ/.
@@ 1 Introduction
Neural Radiance Field (NeRF) [41] has produced state-of-the-art results in novel view synthesis for intricate real-world scenes. NeRF encodes a given scene using Multi-Layer Perceptrons (MLP) and supports queries of density and radiance given 3D coordinates and view directions, which are used to render photo-realistic images from any view points. Moreover, during training, NeRF only requires RGB images with camera poses, which directly links 3D to 2D. The simple but ingenious architecture with its continuous representation quickly starts challenging traditional representations using explicit discrete structures, such as RGB-D images or point clouds. As a result, NeRF is poised to tackle more challenging tasks in 3D vision.
One important downstream task that can benefit from NeRF representations is 3D object segmentation, which is fundamental in 3D vision and widely used in many applications. To address object segmentation in NeRF, researchers have investigated various methods. Semantic-NeRF [65], targeting semantic segmentation, is one of the first works in this direction. DFF [33] distills the knowledge of pre-trained features such as DINO [9] into a 3D feature field for unsupervised object decomposition. Supervised approaches, such as [47], utilize Mask2Former [14] to obtain initial 2D masks and lifts them to 3D with a panoptic radiance field. Although these methods demonstrate impressive results, their performance is constrained by the pre-trained models used to produce features.
Recently, large vision models, such as Segment Anything Model (SAM) [32], have emerged with strong zero-shot generalization performance and can be adopted as the backbone component for many downstream tasks. Specifically, SAM proposes a new paradigm for segmentation tasks which can accept a wide variety of prompts as input, and produce segmentation masks of different semantic levels as output. The versatility and generalizability of SAM thus shed light on a new way to perform promptable object segmentation in NeRF. While there exist some investigations [10, 13, 21] into this area, the mask quality in novel views is still unsatisfactory.
@@
  • We propose SANeRF-HQ, which is one of the first attempts in producing high-quality 3D object segmentation in NeRF in terms of more accurate segmentation boundaries and better multi-view consistency.
  • We validate our method by assembling and evaluating quantitatively on a challenging dataset with high-quality ground-truths.
  • We present a more general framework to embed foundation 2D image models into Neural Radiance Field and extend it to different 3D segmentation tasks in NeRFs.
Comparing with [21] and [10], SANeRF-HQ can produce more accurate segmentation results and is more flexible to a variety of segmentation tasks. Similar to SAM [32] in 2D segmentation, SANeRF-HQ can automatically segment a given entire 3D scene in high quality without any user inputs. SANeRF-HQ inherits the zero-shot performance from SAM, instead of being bounded by pre-trained models with limited generalizability  [38, 5, 47]. SANeRF-HQ is a general framework which can be readily extended to 4D dynamic NeRFs, where temporal consistency can be naturally handled in a similar way as our multi-view consistency. Our preliminary results confirm the potential extension to dynamic scenes.
@@ 2.1 Object Segmentation
2D image segmentation is a thoroughly studied area where a lot of progress has been made, including semantic segmentation [64, 22, 12, 63], instance segmentation [51, 55, 23, 7, 6, 56], and panoptic segmentation [36, 28, 35, 31, 58, 18]. With the emergence of Transformer-based models [8], Vision Transformer (ViT) [19, 39, 60, 61] has become increasingly popular as a backbone structure, and pre-trained features from large models such as MAE [24, 52, 57] have demonstrated great power in segmentation tasks. Moreover, significant research effort [14, 62, 28] focuses on universal models to accomplish multiple segmentation tasks under different training configuration. However, the performance of these models on open-world images is not satisfactory, as they cannot go beyond the limits prescribed by the underlying training datasets.
Thus, latest research has been focusing on open-world segmentation, aiming to generalize segmentation models to unseen data. The recent advancement in visual foundation models have attracted great attention. DINOv2 [44] leverages self-supervised distillation and produces visual features across domains without any fine-tuning. Segment Anything Model (SAM) [32], as a more significant breakthrough, shows promising results on promptable segmentation. Given diverse and plentiful training data, the prompt-based architecture can enable zero-shot generalization, which extends relevant tasks to wider data categories. Recent works [45, 16, 66] have already adopted SAM in many downstream tasks, enhancing the capability of different segmentation models.
@@ 2.2 Segmentation in Neural Radiance Field
The family of Neural Radiance Fields (NeRF) [41, 42, 11, 3, 2, 4] has become the state-of-the-art for novel view synthesis. Beyond that, requiring only multi-view images with camera parameters during training, NeRF can also be regarded as an implicit 3D representation since it captures the 3D structural details of the scenes. Due to the capability of linking 2D images to 3D volumes, NeRF shows potential impact in various 3D visual tasks, and numerous research works have been focusing on 3D object segmentation and scene decomposition in NeRF.
@@
Semantic-NeRF [65] extends NeRF with an additional branch encoding 3D semantic labels for semantic-level segmentation. Other research investigated 3D instance segmentation and make great efforts in solving inconsistency across views. For example, [38] adopts an 3D object detection method [26] to resolve 2D mask correspondences. [47] utilize linear assignment, while [5] employs a contrastive loss to optimize 3D instance embeddings. Besides, there are unsupervised methods [48, 59, 43], which can separate foreground objects from background. However, these approaches are mostly limited to simple scenes and often struggle to generalize to complex open-world problems. To fully utilize 2D features from pre-trained models, some research [30, 21, 33] introduces an extra feature field to the vanilla NeRF model, which can fuse 2D features from pre-trained models into 3D. For instance, LERF [30] splits images into patches of different sizes to obtain multi-scale CLIP [27] feature maps that can supervise the neural field training. ISRF [21] uses DINO [9] features and K-Means clustering to separate user-selected regions from background. However, lacking a powerful decoder, the segmentation based on these features is not accurate and sharp along the boundaries.
With the proposal of SAM, SA3D [10] employs SAM on the NeRF-rendered images and achieve 3D segmentation from a single-view 2D mask by self-prompting. Nonetheless this pipeline relies on the first-view mask and is susceptible to the ambiguity inherent in SAM in delineating intricate structures during its self-prompting. SAN [13] chooses to distill the SAM encoder with a neural field to render SAM feature maps from novel views, which are supplied to the SAM decoder to produce the segmentation. However, it can produce inconsistent masks in different views, and distilling low resolution SAM feature maps results in aliasing in the output masks, in the form of jagged mask edges. Contrary to these approaches, our proposed method directly aggregates masks for neural field training, which naturally addresses the aliasing and consistency issues.
@@ 3 Method
Given a pre-trained NeRF, our method aims to segment any kind of object in 3D, conditioning on the manual prompts and/or other user supplied inputs. The SANeRF-HQ pipeline in Figure 1 consists of three major components: a feature container, a mask decoder, and a mask aggregator. The feature container encodes the SAM features of images. The mask decoder propagates the user supplied hints between different views and generates intermediate mask outputs using the SAM features from the container. Finally, the mask aggregator integrates the resulted 2D masks into 3D space and utilizes the color and density fields from NeRF models to achieve high-quality 3D segmentation.
@@ 3.1 Feature Container
The first step of utilizing SAM is to encode the images into 2D features using the SAM feature encoder. These features can be used repeatedly when predicting and propagating masks, thus they can be precomputed or distilled for a scene and reused for different input prompts.
We consider two different methods for the feature container. The first method is to compute and cache the features of multiple views. This allows us to reuse ground-truth SAM features for different user prompts and generate accurate 2D mask when decoding. However, the cache size is constrained by the memory available. It also requires extra time to run the encoder if users choose to supply prompts on any of the uncached novel views.
@@
The feature field enables feature map rendering from any viewpoint efficiently, as aggregating features in the neural field is typically faster than running the original SAM encoder. However, the feature maps produced by the SAN encoder has a relatively low resolution, which can cause severe aliasing in the rendered feature maps. While this can be alleviated by augmenting the input camera views and sample more rays during distillation, the rendered features still deteriorate after distillation. The rendered SAN features usually fail to retain accurate high frequency spatial information along boundaries, which consequently leads to jagged mask boundaries after decoding.
Noting the complementary advantages and disadvantages of the caching method and the feature distillation method, we conducted experiments on both methods in the ablation study.
@@ 3.2 Mask Decoder
where $\hat{\Phi}$ is the feature map. NeRF can estimate depth with Equation 5,
@@ 3.3.1 Object Field
Given the decoder output from different views, the object field can fuse all 2D images and generate accurate 3D masks. The mask is represented by a $L$-dimensional object identity vector $i$. To represent the identity value, an additional branch is introduced into the pre-trained NeRF model, parameterized as $\Theta_{o}$. Different from the function of RGB color, which is view-dependent, object identity function is defined by $\mathbf{i}=f(\mathbf{x};\Theta_{o})$, where the view direction vector is not included in the inputs due to the view invariance of the object masks in 3D. The equation for mask rendering is similar to that of image rendering:
@@
where $\sigma(t_{k})$ is inherited from the pre-trained NeRF model. Volume density $\sigma$ which interprets the 3D geometry in NeRF makes the object field aware of the structural information. The object field is trained with the cross-entropy loss $\mathcal{L}_{o}$:
@@ 4 Experiments
Figure 4: Comparison with SA3D and ISRF on the Table. SANeRF-HQ can preserve structure details of the table.
@@ 4.1 Datasets
We evaluate our performance on data from multiple datasets, mixed with synthetic and real-world scenes. We compare the masks projected on 2D images with ground-truth masks to evaluate our results quantitatively. For data without ground-truth, we manually annotate object masks. We categorize the dataset used into the following 5 groups:
  • Mip-NeRF 360: a dataset widely used in NeRF research including synthetic and real-world examples. In our experiments, we use the data in [3].
  • LERF: a set of scenes captured by [30], which contains complex real-world samples.
  • LLFF: first used in [40], the dataset contains scenes with only front views. We use the masks released with [21].
@@
For each group above, we select scenes with representative objects and segmentation outcomes that can be clearly identified by humans. In total, 24 scenes are selected, each containing 1 to 3 object segmentation. For those without ground-truth masks, we use SAM and CascadePSP [15] with manual annotation to create ground-truth.
@@
4.2 Metrics
We evaluate the segmentation performance of SANeRF-HQ using mean Intersection over Union (mIoU) and Accuary (Acc). For each object, we render the object masks in several novel test views, then combine the pixels in all test views to calculate the per-object IoU and Acc. The overall mIoU and Acc are averaged over all objects.
4.3 Comparison
We provide the comparison with some zero-shot segmentation methods mentioned in Section 2.2. These methods leverages large vision models such as SAM [32] or DINO [9] and can achieve zero-shot segmentation on general scenes when user prompts are provided. Table 1 shows the quantitative comparison with 4 methods on 5 different datasets mentioned in Section 4.1. We use point prompts as they are more unequivocal.
To ensure fair comparison and eliminate bias in prompting, we use the same point prompts to get the initial masks for SA3D [10]. In addition, since SA3D requires a single-view mask as input, we manually select a view containing the major component of the target object, and pick the mask that best matches the ground-truth, instead of using the predicted scores from SAM to select the initial mask. To our best extent, we make sure that SA3D is provided with good initialization. For ISRF [32], strokes are used as prompts, so we manually connect the point prompts to create strokes. Figure 3 and  4 demonstrate the qualitative comparison. SA3D uses a self-prompting strategy and iteratively inverse renders the 2D masks to a voxel grid, whereas our method uses a set of global prompts and collectively optimize the object field. Despite the use of IoU rejection, self-prompting may incorrectly include occluded regions in novel views into prompts, which can accumulate errors, especially in initial iterations. ISRF lifts DINO features [9] into a neural field, but its clustering and searching process produces less accurate mask boundaries compared to SAM.
Figure 6: Qualitative Results of text prompts and auto segmentation.
4.4 More Qualitative Results
Figure 6 illustrates the qualitative results of our methods on other segmentation tasks. Utilizing Grounding-DINO [37], our method is also capable of segmenting objects based on text prompts. Additionally, our method can perform automatic 3D segmentation, utilizing the video rendered by NeRF, the auto-segmentation function of SAM, and incorporating [16] into the mask decoder. These results show the potential of our method in readily extending to dynamic NeRF.
4.5 Ablation Study

### 06 Apr 2024 v2

@@ Abstract
Recently, the Segment Anything Model (SAM) has showcased remarkable capabilities of zero-shot segmentation, while NeRF (Neural Radiance Fields) has gained popularity as a method for various 3D problems beyond novel view synthesis. Though there exist initial attempts to incorporate these two methods into 3D segmentation, they face the challenge of accurately and consistently segmenting objects in complex scenarios. In this paper, we introduce the Segment Anything for NeRF in High Quality (SANeRF-HQ) to achieve high-quality 3D segmentation of any target object in a given scene. SANeRF-HQ utilizes SAM for open-world object segmentation guided by user-supplied prompts, while leveraging NeRF to aggregate information from different viewpoints. To overcome the aforementioned challenges, we employ density field and RGB similarity to enhance the accuracy of segmentation boundary during the aggregation. Emphasizing on segmentation accuracy, we evaluate our method on multiple NeRF datasets where high-quality ground-truths are available or manually annotated. SANeRF-HQ shows a significant quality improvement over state-of-the-art methods in NeRF object segmentation, provides higher flexibility for object localization, and enables more consistent object segmentation across multiple views. Results and code are available at the project site: https://lyclyc52.github.io/SANeRF-HQ/.
22footnotetext: This research is supported in part by the Research Grant Council of the Hong Kong SAR under grant no. 16201420.
@@ 1 Introduction
Neural Radiance Field (NeRF) [41] has produced state-of-the-art results in novel view synthesis for intricate real-world scenes. NeRF encodes a given scene using Multi-Layer Perceptrons (MLPs) and supports queries of density and radiance given 3D coordinates and view directions, which are used to render photo-realistic images from any view points. Moreover, during training, NeRF only requires RGB images with camera poses, which directly links 3D to 2D. The simple but ingenious architecture with its continuous representation has quickly started to challenge traditional representations using explicit discrete structures, such as RGB-D images or point clouds. As a result, NeRF is poised to tackle more challenging tasks in 3D vision.
One important downstream task that can benefit from NeRF representations is 3D object segmentation, which is fundamental in 3D vision and widely used in many applications. Researchers have investigated applying NeRF to object segmentation. Semantic-NeRF [66], targeting semantic segmentation, is one of the first works in this direction. DFF [33] distills the knowledge of pre-trained features such as DINO [9] into a 3D feature field for unsupervised object decomposition. Supervised approaches, such as [47], utilize Mask2Former [15] to obtain initial 2D masks and lifts them to 3D with a panoptic radiance field. Although these methods demonstrate impressive results, their performance is constrained by the pre-trained models used to produce features.
Recently, large vision models such as Segment Anything Model (SAM) [32] have emerged. These models with strong zero-shot generalization performance can be adopted as the backbone component for many downstream tasks. SAM presents a new paradigm for segmentation tasks. It can accept a wide variety of prompts as input, and produce segmentation masks of different semantic levels as output. The versatility and generalizability of SAM suggest a new way to perform promptable object segmentation in NeRF. While there exist some investigations [10, 14, 22] into this area, the mask quality in novel views is still unsatisfactory.
@@
  • We propose SANeRF-HQ, one of the first attempts at producing high-quality 3D object segmentation in NeRF in terms of more accurate segmentation boundaries and better multi-view consistency.
  • We validate our method by evaluating quantitatively on a challenging dataset with high-quality ground-truths.
  • We present a general framework to embed foundation 2D image models into NeRFs and extend it to different 3D segmentation tasks in NeRFs.
Comparing with [22] and [10], SANeRF-HQ can produce more accurate segmentation results and is more flexible to a variety of segmentation tasks. SANeRF-HQ inherits the zero-shot performance from SAM, instead of being bounded by pre-trained models with limited generalizability  [38, 5, 47]. Moreover, SANeRF-HQ is a general framework, which has the hidden capability to automatically segment a given 3D scene like SAM [32] in 2D segmentation. Also, SANeRF-HQ can be potentially extended to 4D dynamic NeRFs, where temporal consistency can be handled in a similar way as our multi-view consistency. Our preliminary results demonstrate promising prospect for the extension to dynamic scenes.
@@ 2.1 Object Segmentation
2D image segmentation including semantic segmentation [65, 23, 13, 64], instance segmentation [52, 56, 24, 7, 6, 57], and panoptic segmentation [36, 28, 35, 31, 59, 19] is a thoroughly studied area. With the emergence of Transformer-based models [8], Vision Transformer (ViT) [20, 39, 61, 62] has become increasingly popular as a backbone structure, and pre-trained features from large models such as MAE [25, 53, 58] have demonstrated great power in segmentation tasks. Meanwhile, significant research effort [15, 63, 28] has focused on universal models to accomplish multiple segmentation tasks under different training configurations. However, the performance of these models on open-world images is not satisfactory, as they cannot go beyond the limits prescribed by the underlying training datasets.
In response, the latest research has been focusing on open-world segmentation, aiming to generalize segmentation models to unseen data. The recent advancement in visual foundation models have attracted great attention. DINOv2 [43] leverages self-supervised distillation and produces visual features across domains without any fine-tuning. SAM [32], a more significant breakthrough, shows promising results on promptable segmentation. Given diverse and plentiful training data, the prompt-based architecture can enable zero-shot generalization, which extends relevant tasks to wider data categories. Recent works [45, 17, 67] have already adopted SAM in many downstream tasks, enhancing the capability of different segmentation models.
@@ 2.2 Segmentation in Neural Radiance Field
NeRF [41, 51, 11, 3, 2, 4] has become the state-of-the-art for novel view synthesis. Beyond that, requiring only multi-view images with camera parameters during training, NeRF can also be regarded as an implicit 3D representation since it captures the 3D structural details of the scenes. Due to its capability of linking 2D images to 3D volumes, NeRF shows potential impact in various 3D visual tasks, and numerous research works have been focusing on 3D object segmentation and scene decomposition in NeRF.
@@
[66, 12] extend NeRF with an additional branch encoding 3D semantic labels for semantic-level segmentation. Other research investigates 3D instance segmentation and makes great efforts in solving inconsistency across views. For example, [38] adopts an 3D object detection method [27] to resolve 2D mask correspondences. [47] utilizes linear assignment, while [5] employs a contrastive loss to optimize 3D instance embeddings. There are also unsupervised methods [48, 60, 42], which can separate foreground objects from background. However, all aforementioned approaches are mostly limited to simple scenes and often struggle to generalize to complex open-world problems. To fully utilize 2D features from pre-trained models, some research [30, 22, 33] introduces an extra feature field to the vanilla NeRF model, which can fuse 2D features from pre-trained models into 3D. For instance, LERF [30] splits images into patches of different sizes to obtain multi-scale CLIP [44] feature maps that can supervise the neural field training. ISRF [22] uses DINO [9] features and K-Means clustering to separate user-selected regions from background. However, lacking a powerful decoder, the segmentation based on these features is not accurate and sharp along the boundaries.
With the proposal of SAM, SA3D [10] employs SAM on the NeRF-rendered images and achieves 3D segmentation from a single-view 2D mask by self-prompting. Nonetheless this pipeline relies on the first-view mask and is susceptible to the ambiguity inherent in SAM in delineating intricate structures during its self-prompting. SAN [14] chooses to distill the SAM encoder with a neural field to render SAM feature m aps from novel views, which are supplied to the SAM decoder to produce the segmentation. However, it can produce inconsistent masks in different views, and distilling low-resolution SAM feature maps results in aliasing in the output masks, in the form of jagged mask edges. Contrary to these approaches, our proposed method directly aggregates masks for neural field training, which naturally addresses the aliasing and consistency issues.
@@ 3 Method
Given a pre-trained NeRF, our method aims to segment any target object in 3D, conditional on the manual prompts and/or other user supplied inputs. The SANeRF-HQ pipeline, shown in Figure 1, consists of three major components: a feature container, a mask decoder, and a mask aggregator. The feature container encodes the SAM features of images. The mask decoder propagates the user-supplied prompts between different views and generates intermediate mask outputs using the SAM features from the container. Finally, the mask aggregator integrates the resultant 2D masks into 3D space and utilizes the color and density fields from NeRF models to achieve high-quality 3D segmentation.
@@ 3.1 Feature Container
The first step of utilizing SAM is to encode the images into 2D features using the SAM encoder. These features can be used repeatedly when predicting and propagating masks, thus can be pre-computed or distilled for a scene and reused for different prompts.
We consider two methods for the feature container. The first method is to compute and cache the features of multiple views. This allows us to reuse ground-truth SAM features for different user prompts and generate accurate 2D masks when decoding. However, the cache size is constrained by the memory available. This method also requires extra time to run the encoder if users choose to supply prompts on any of the uncached novel views.
@@
The feature field enables efficient feature map rendering from any viewpoint , as aggregating features in the neural field is typically faster than running the original SAM encoder. However, the feature maps produced by the SAN encoder have a relatively low resolution, which can cause severe aliasing in the rendered feature maps. While this can be alleviated by augmenting the input camera views and sampling more rays during distillation, the rendered features still deteriorate after distillation. The rendered SAN features usually fail to retain accurate high-frequency spatial information along boundaries, which consequently leads to jagged mask boundaries after decoding.
Noting the complementary advantages and disadvantages of the caching method and feature distillation method, we conducted experiments on both in the ablation study.
@@ 3.2 Mask Decoder
where $\hat{\Phi}$ is the feature map. NeRF can estimate depth with
@@ 3.3.1 Object Field
Given the decoder output from different views, the object field can fuse all 2D images and generate accurate 3D masks. The mask is represented by an $L$-dimensional object identity vector $i$. To represent the identity value, an additional branch is introduced into the pre-trained NeRF model, parameterized as $\Theta_{o}$. Different from the function of RGB color, which is view-dependent, the object identity function is defined by $\mathbf{i}=f(\mathbf{x};\Theta_{o})$, where the view direction vector is not included in the inputs due to the view invariance of the object masks in 3D. The equation for mask rendering is similar to that of image rendering:
@@
where $\sigma(t_{k})$ is inherited from the pre-trained NeRF model. The volume density $\sigma$, which interprets the 3D geometry in NeRF, makes the object field aware of the structural information. The object field is trained with the cross-entropy loss $\mathcal{L}_{o}$:
@@ 4 Experiments
Figure 4: Comparison with SA3D and ISRF on the Garden. SANeRF-HQ can preserve structure details of the table.
@@ 4.1 Datasets
We evaluate our performance on data from multiple datasets, containing synthetic and real-world scenes. We compare the masks projected on 2D images with ground-truth masks to evaluate our results quantitatively. For data without ground-truth, we manually annotate object masks. We use five datasets:
  • Mip-NeRF 360: a dataset widely used in NeRF research that includes synthetic and real-world examples. In our experiments, we use the data from [3].
  • LERF: a set of scenes captured in [30], which contains complex real-world samples.
  • LLFF: first used in [40], the dataset contains scenes with only front views. We use the masks released in [22].
@@
For each group above, we select scenes with representative objects and segmentation outcomes that can be clearly identified by humans. In total, 24 scenes are selected, each containing 1 to 3 object segmentations. For those without ground-truth masks, we use SAM and CascadePSP [16] with manual annotation to create the ground-truth.
4.2 Metrics
We use mean Intersection over Union (mIoU) and Accuary (Acc) as the evaluation metrics. For each object, masks in novel test views are rendered and results from all test views are combined to compute the per-object IoU and Acc. The overall mIoU and Acc are averaged over all objects.
4.3 Comparison
We provide the comparison with some zero-shot segmentation methods mentioned in Section 2.2. These methods leverages large vision models such as SAM [32] or DINO [9], and can achieve zero-shot segmentation on general scenes given user prompts. Table 1 shows the quantitative comparison with four methods on 5 datasets listed in Section 4.1. We use point prompts as they are less equivocal. For fair comparison, the same point prompts are used to get the initial masks for SA3D [10]. As SA3D requires a single-view mask as input, we manually select a view containing the major component of the target object and pick the mask that best matches the ground-truth, instead of using the predicted scores from SAM. To our best extent, we ensure that SA3D is provided with good initialization. ISRF [22] uses strokes as prompts, hence we manually connect multiple point prompts to create strokes. Figure 3 and  4 demonstrate the qualitative comparison. Table 1 presents the quantitative results. Our method outperforms others in the comparison quantitatively on all datasets. SA3D uses a self-prompting strategy and iteratively inverse renders the 2D masks to a voxel grid, whereas our method uses a set of global prompts and collectively optimizes the object field. Despite the use of IoU rejection, self-prompting may incorrectly include occluded regions in novel views into prompts, which can accumulate errors in predicted masks, especially in initial iterations. This sensitivity to inaccurate SAM predictions partially explains SA3D’s underperformance on the LERF and 3D-FRONT dataset, as the former contains scenes with small, partially-occluded objects, while the latter has furniture that may introduce semantic ambiguity. ISRF lifts DINO features [9] into a neural field, but its clustering and searching process produces less accurate mask boundaries compared to SAM.
@@ 4.3 Comparison
Backbone Container Mip-NeRF 360 LERF LLFF 3D-FRONT Others
SAM Cache 99.3 93.6 98.9 90.3 99.5 95.8 97.6 84.8 99.6 91.1
Distillation 99.2 91.0 99.0 90.7 99.3 95.2 98.6 89.9 99.6 91.1
SAM-HQ Cache 99.4 94.4 99.2 93.1 99.6 96.8 98.9 91.8 99.5 90.1
Table 2: Quantitative Results on Different Backbones and Feature Containers.
4.4 Ablation Studies
SAM Models and Feature Containers. In Section 3.1, two methods to store the SAM features are proposed. We investigate the performance of these two methods, as well as two pre-trained segmentation models, original SAM and HQ-SAM [29], on our evaluation set. Table 2 demonstrates the quantitative results. HQ-SAM with a cache container performs the best in most cases, whereas HQ-SAM with distillation is not included due to heavier overhead of distilling multi-level feature maps. We find no significant performance gap between feature distillation and caching over different datasets, hence the difference in computational cost may play a more important role, which we discuss in the supplementry material.